A sample of Butane gas was found in a victim’s lungs after an autopsy. Butane is an organic compound with the chemical formula C4H10

If a sample of Butane contains 2.58 x 10²⁸ Carbon atoms, how many atoms of hydrogen are present?

If a sample of Butane contains 3.25 x 10³⁴ Hydrogen atoms, how many Carbon atoms are present?
1
Expert's answer
2019-12-03T08:24:26-0500

(i)


"N(C) : N(H) = 4 : 10 = 1 : 2.5"

"N(H) = N(C)*2.5 = 2.58*10^{28}*2.5 = 6.45*10^{28}"

(ii)


"N(C) = {3.25*10^{28} \over 2.5} = 1.3 * 10^{28}"